According to a report from the research for Studying Health System Change, 20% of South Africans delay or go without medical care because of concerns about cost. Suppose that 8 individuals are randomly selected. (a) What is the probability that two individuals will delay or go without medical care? (2) (b) What is the probability that at most two individuals will delay or go without medical care? (3) (c) What is the probability that at least seven individuals will delay or go without medical care? (3)


1
Expert's answer
2022-06-14T13:08:11-0400

Let "X=" the number of individuals who will delay or go without medical care: "X\sim Bin(n, p)."

Given "n=8, p=0.2, q=1-p=0.8"

a)


"P(X=2)=\dbinom{8}{2}(0.2)^{2}(0.8)^{8-2}""=0.29360128"


b)


"P(X\le2)=P(X=0)+P(X=1)+P(X=2)""=\dbinom{8}{0}(0.2)^{0}(0.8)^{8-0}+\dbinom{8}{1}(0.2)^{1}(0.8)^{8-1}"

"+\dbinom{8}{2}(0.2)^{2}(0.8)^{8-2}=0.16777216"

"+0.33554432+0.29360128"


"=0.79691776"



c)


"P(X\ge 7)=P(X=7)+P(X=8)"

"=\dbinom{8}{7}(0.2)^{7}(0.8)^{8-7}+\dbinom{8}{8}(0.2)^{8}(0.8)^{8-8}"

"=0.00008192+0.00000256"

"=0.00008448"